# How does online shopping affect offline price sensitivity?

### v1

@@ 6.4 Demand Model Parameter Estimates
Table 6 presents the estimation results of the demand model based on Equation (7). The standard errors of the parameter estimates are derived using block bootstrap with 500 replications. All the price coefficients are negative and statistically significant. The price interaction terms are also negative and significant for dry dog food and for dog and cat hygiene categories but are insignificant for dry cat food. We discuss the potential sources for this heterogeneity in §$\S$7.1.
The coefficient of the interaction term in the logit model should not be interpreted directly. Instead, the correct way to interpret the interaction effect is to calculate the marginal effect and price elasticity (Ai and Norton 2003; Karaca-Mandic et al. 2012). Accordingly, we calculate price elasticities using the statistically significant coefficients, as specified in Equation (14) and discuss their implications in the next section.
@@ 7 ATT of Online Adoption on Offline Price Elasticity
Now that we have the demand parameter estimates, we then measure the category-level own-price elasticity and formalize the ATT of online adoption on offline price elasticity. In §$\S$7.1, we present the elasticity estimates from the estimation procedure. Next, in §$\S$7.2, we perform a within-customer comparison of adopters’ offline price elasticity before and after the adoption of online shopping. In §$\S$7.3, we formalize the changes in offline price elasticity due to the adoption of online shopping using a staggered difference-in-differences (DID) model.
@@ 7.1 Elasticity Estimates
$\quad\eta_{ijt}=\frac{\partial Pr_{ijt}}{\partial\textrm{Price}_{jt}}\cdot\frac{\textrm{Price}_{jt}}{Pr_{ijt}}=(\alpha_{0}+\alpha_{2}\times\textrm{cumPercOnlineSpend}_{it})\times\textrm{Price}_{jt}\times(1-Pr_{ijt})$ (14)
@@
Before online adoption, consumers’ offline price elasticities closely mirror those of non-adopters, suggesting comparable pre-trends. However, after adopting online shopping, adopters consistently become more price sensitive in their offline shopping. The magnitude of the increase in offline price elasticity is most pronounced for dog hygiene and cat hygiene, moderate for dry dog food, and barely visible for dry cat food. The sharp increases in elasticities for the hygiene categories are likely due of the fact that these products have low switching costs: owners face little risk in trying a new hygienic mat for their dog if it is lower priced, with other attributes like pad size being standardized. The smaller changes to price elasticities in the two pet food categories, and a zero effect for cat food, is likely due to the fact that brand switching is not so simple in the case of pet food. Changing a dog’s or cat’s diet entails palatability trials, digestive risks, and rejection. Cats in particular, are very picky eaters and have especially strong food preferences in terms of shape, smell and texture (Nichelason 2024). Since the main coefficient of interest in Table 6 for cat food is not statistically significant at the 5% level, and the elasticity figure also shows no effect for cat food, we drop this category in our subsequent analyses.
@@ 7.2 Within-Customer Analysis
$\mathcal{I}\{t\geq T_{i}^{\textrm{adopt}}\}$0.3870∗∗∗ -0.1173 -1.916∗∗∗ -0.7239∗∗ -0.8860∗∗∗ -0.3347∗∗
@@ 7.3 Difference-In-Differences (DID) Analysis
Although TWFE regression is widely used in empirical studies involving staggered adoption, it only provides consistent estimates of ATT under the assumption of homogeneous treatment effects, a condition that is rarely met in practice; see De Chaisemartin and d’Haultfoeuille (2020); Callaway and Sant’Anna (2021); Sun and Abraham (2021); Goodman-Bacon (2021) for detailed discussions of this issue. Specifically, Goodman-Bacon (2021) shows that the TWFE estimator $(\phi_{TWFE})$ effectively represents a weighted average of all possible 2-by-2 combinations of pairwise difference-in-differences comparisons between groups of units treated at different times.
To address this issue, we adopt the staggered difference-in-differences (DID) framework proposed by Callaway and Sant’Anna (2021), hereafter referred to as the CS estimator as our main specification. The CS estimator addresses the biases associated with TWFE estimators by eliminating problematic comparisons between late-treated and already-treated units and has been applied recently in the marketing and economics literature (Ang 2021; Braghieri et al. 2022; Bekkerman et al. 2023). This method is particularly suited to our context, because we have customers adopting online shopping over a one year period, in 12 cohorts (one cohort per month). As discussed earlier, because there were exogenous events during this time period (e.g., the onset of COVID-19), it is possible that the ATT is different across cohorts. Therefore, we use this estimator in our main analysis and then supplement the analysis with TWFE estimates as a robustness check (which provide similar results in terms of direction and magnitude). We refer the reader to the original paper for a discussion of the technical details of how the CS estimator works. The main idea is to identify the Average Treatment Effect on the Treated (ATT) for each post-treatment period, for consumers who are treated at different times, and then calculate the overall treatment effect as a weighted average of these ATTs. In our context, the ATT in month $t$ for customers who adopt online shopping in month $g$ can be defined as:
@@
As discussed by Callaway and Sant’Anna (2021), this effect, $ATT(g,t)$, can be identified provided that parallel pre-trends hold. We provide a test of parallel pre-trends in Figure A7 in Web Appendix §E.3.
@@
We present the overall treatment effect estimates $(\theta_{o})$ in Table 8. Since the customer-month level price elasticities are themselves estimates from our demand model with confidence intervals around them (and not directly observed from data), we compute standard errors for our ATT estimates using customer-level block bootstraps with 500 replications. Figure A3 in Web Appendix §D shows the bootstrapped distribution of these ATT estimates. The results in Table 8 suggest that consumers become more price-sensitive in the offline channel after they adopt online shopping, and this finding is consistent across these categories. The effect is stronger for the dog and cat hygiene categories, consistent with the elasticity visualization in Figure 2. To help with interpretation, we calculate the relative magnitude of the elasticity increase in percentage terms by dividing the ATT by the average pre-adoption price elasticity of adopters and show these percentage changes in the second row of Table 8. The most substantial effects are observed for dog hygiene and cat hygiene categories; offline price elasticity increases by 72.7% (i.e., -1.792 / -2.464) for dog hygiene and 34.7% (i.e., -0.738 / -2.125) for cat hygiene. The increase in offline price elasticity is lower for dry dog food, at 8.4% (i.e., -0.248 / -2.935). As discussed in §7.1, the pronounced increase in offline price elasticity for hygiene products (vs. dry food products) is likely driven by the lower switching costs associated with trying new brands of hygiene mats or other hygiene-related items. In summary, consistent with the descriptive evidence earlier, we find that consumers become more price-sensitive in their offline shopping trips when they start shopping online.
@@ 8.1 Generalizability
Second, we examine the extent to which our results are likely to be driven by the fact that online prices are slightly lower than offline prices in our setting. To that end, we examine the magnitude of the price differences between online and offline channels in our study (see details in in Web Appendix E.1). We find that online prices are only 2–6% lower than offline in our data. Furthermore, the variation in prices explained by channel is significantly smaller than the variation in prices explained by SKU-level differences – across all categories, only 4% of the variation in prices is explained by channels, whereas when we include SKU fixed effects, this increases to 65%. Overall, these data patterns suggest that the differences between online and offline prices are relatively small and are unlikely to be the primary driver of our results. Other potential mechanisms, such as consumers becoming more accustomed to comparing prices across brands or their exposure to a wider variety of prices, are likely to be at play as well (as discussed in §$\S$1). Thus, we expect that our results are likely to generalize to other empirical settings where online prices are the same or higher than offline prices. [^9] In practice, there is considerable heterogeneity in whether online prices are lower, the same, or higher than offline prices across retailers, countries, and product categories (Cavallo 2017).
@@ 8.2 Robustness checks
We first examine alternative demand specifications in §$\S$8.2.1 and then consider matching methods and alternative control groups in §$\S$8.2.2 and 8.2.3, respectively.
@@ 8.2.1 Alternative Demand Specifications
    Accounting for a potential change in state dependence or brand loyalty due to online adoption: In our main model, we include customer-brand fixed effects to capture time-invariant preferences and use $\mathcal{I}\{\textrm{BuyLastMonth}_{jt}\}$ to capture the state dependence. We now consider a specification with an additional control variable, $\mathcal{I}\{\textrm{BuyLastMonth}_{jt}\}\times\textrm{cumPercOnlineSpend}_{it}$, which captures the interaction between state dependence (e.g., repeated purchase of a brand) and the share of online spending. This can account for any effects that online adoption may have on brand loyalty.
The utility specifications and the estimation results from these two exercises are shown in Web Appendix §$\S$§E.2. We find that the results are consistent with our main specification. We also do not find evidence of significant changes to offline brand loyalty after consumers adopt online shopping.
@@ 8.2.2 Matching and IPW Approaches to Control for Selection on Observables
Consumers self-select into adopting a retailer’s online channel (or not), and as shown in Web Appendix §$\S$§A.2, there are some systematic differences between adopters and non-adopters in terms of observable characteristics like demographics. In general, such differences are not an issue as long as the parallel pre-trends assumption holds (see Chapter 5 of Angrist and Pischke (2008) and Roth et al. (2023) for additional discussions). To that end, we consider the pre-adoption trends in price elasticities between adopters and non-adopters and verify that the parallel pre-trends assumption holds in Web Appendix §$\S$§E.3.
@@
We perform the propensity score estimation and matching for each adoption cohort separately. Specifically, we define the pre-adoption period separately for each cohort and measure pre-adoption variables (e.g., monthly average spend) using all purchase data up to the month of adoption (Gu and Kannan 2021). For example, when identifying matched non-adopters for customers who adopt in July 2019, we measure the pre-adoption variables for both adopters and non-adopters using the data up to June 30, 2019, and estimate the propensity score model for the adoption cohort in July 2019. We repeat this process for each cohort and estimate twelve different propensity score models corresponding to each monthly adoption cohort. We compute propensity scores with a logistic regression model and the variables included are demographics (e.g., gender, age, and household income) and pre-adoption purchase variables (i.e. average monthly spend, number of unique orders, unique brands, and unique categories). With the estimated propensity scores, we use our CS estimator to estimate the ATT using both IPW and PSM methods. The propensity score model specification and detailed results from the IPW and PSM methods are shown in Web Appendix E.4. We find that our conclusions remain the same.

### v2

@@ 6.4 Demand Model Parameter Estimates
Table 6 presents the estimation results of the demand model based on Equation (7). The standard errors of the parameter estimates are derived using block bootstrap with 500 replications. All the price coefficients are negative and statistically significant. The price interaction terms are negative and significant for three out of the four categories we study: dry dog food and for dog and cat hygiene categories, indicating that the adoption (and extent of adoption) of a retailer’s channels makes consumers more price sensitive during their subsequent offline shopping trips at that retailer’s physical stores. The coefficient is insignificant for dry cat food. We discuss the potential sources for this heterogeneity in $\S$7.1.
The magnitude of the coefficient of the interaction term in the logit model is difficult to interpret directly. Instead, we can interpret the magnitude of the effect we find by computing the marginal effects and price elasticities (Ai and Norton, 2003; Karaca-Mandic et al., 2012). To that end, we calculate price elasticities using the statistically significant coefficients, as specified in Equation (14) and discuss their implications in the next section.
@@ 7 ATT of Online Adoption on Offline Price Elasticity
Now that we have the demand parameter estimates and we see a negative and significant effect on the interaction term of interest, we measure the category-level own-price elasticities and formalize the ATT of the adoption of a retailer’s online channels on offline price elasticities. In $\S$7.1, we present the elasticity estimates from the estimation procedure. Next, in $\S$7.2, we perform a within-customer comparison of adopters’ offline price elasticity before and after the adoption of online shopping. In $\S$7.3, we formalize the changes in offline price elasticity due to the adoption of online shopping using a staggered difference-in-differences (DID) model.
@@ 7.1 Elasticity Estimates
$\hskip-36.135pt\begin{aligned} \quad\eta_{ijt}=\frac{\partial Pr_{ijt}}{\partial\textrm{Price}_{jt}}\cdot\frac{\textrm{Price}_{jt}}{Pr_{ijt}}=(\alpha_{0}+\alpha_{2}\times\textrm{cumPercOnlineSpend}_{it})\times\textrm{Price}_{jt}\times(1-Pr_{ijt})\end{aligned}$ (14)
@@
Before online adoption, consumers’ offline price elasticities closely mirror those of non-adopters, suggesting comparable pre-trends. However, after adopting online shopping, adopters consistently become more price sensitive in their offline shopping. The magnitude of the increase in offline price elasticity is most pronounced for dog hygiene and cat hygiene, moderate for dry dog food, and barely visible for dry cat food. The sharp increases in elasticities for the hygiene categories are likely due of the fact that these products have low switching costs: owners face little risk in trying a new hygienic mat for their dog if it is lower priced, with other attributes like pad size being standardized. The smaller changes to price elasticities in the two pet food categories, and a zero effect for cat food, are likely due to the fact that brand switching is not so simple in the case of pet food. Changing a dog’s or cat’s diet entails palatability trials, digestive risks, and rejection. Cats, in particular, are very picky eaters and have especially strong food preferences in terms of shape, smell, and texture (Nichelason, 2024). Since the main coefficient of interest in Table 6 for cat food is not statistically significant at the 5% level, and the elasticity figure also shows no effect for cat food, we drop this category in our subsequent analyses.
@@ 7.2 Within-Customer Analysis
$\mathcal{I}\{t\geq T_{i}^{\textrm{adopt}}\}$ -0.3870∗∗∗ -0.1173 -1.916∗∗∗ -0.7239∗∗ -0.8860∗∗∗ -0.3347∗∗
@@ 7.3 Difference-In-Differences (DID) Analysis
Although TWFE regression is widely used in empirical studies involving staggered adoption, it only provides consistent estimates of ATT under the assumption of homogeneous treatment effects, a condition that is rarely met in practice; see De Chaisemartin and d’Haultfoeuille, (2020); Callaway and Sant’Anna, (2021); Sun and Abraham, (2021); Goodman-Bacon, (2021) for detailed discussions of this issue. Specifically, Goodman-Bacon, (2021) shows that the TWFE estimator $(\phi_{TWFE})$ effectively represents a weighted average of all possible 2-by-2 combinations of pairwise difference-in-differences comparisons between groups of units treated at different times.
To address this issue, we adopt the staggered difference-in-differences (DID) framework proposed by Callaway and Sant’Anna, (2021), hereafter referred to as the CS estimator as our main specification. The CS estimator addresses the biases associated with TWFE estimators by eliminating problematic comparisons between late-treated and already-treated units and has been applied recently in the marketing and economics literature (Ang, 2021; Braghieri et al., 2022; Bekkerman et al., 2023). This method is particularly suited to our context, because we have customers adopting online shopping over a one year period, in 12 cohorts (one cohort per month). As discussed earlier, because there were exogenous events during this time period (e.g., the onset of COVID-19), it is possible that the ATT is different across cohorts. Therefore, we use this estimator in our main analysis and then supplement the analysis with TWFE estimates as a robustness check (which provide similar results in terms of direction and magnitude). We refer the reader to the original paper for a discussion of the technical details of how the CS estimator works. The main idea is to identify the Average Treatment Effect on the Treated (ATT) for each post-treatment period, for consumers who are treated at different times, and then calculate the overall treatment effect as a weighted average of these ATTs. In our context, the ATT in month $t$ for customers who adopt online shopping in month $g$ can be defined as:
@@
As discussed by Callaway and Sant’Anna, (2021), this effect, $ATT(g,t)$, can be identified provided that parallel pre-trends hold. We provide a test of parallel pre-trends in Figure A7 in Web Appendix §E.3.
@@
We present the overall treatment effect estimates $(\theta_{o})$ in Table 8. Since the customer-month level price elasticities are themselves estimates from our demand model with confidence intervals around them (and not directly observed from data), we compute standard errors for our ATT estimates using customer-level block bootstraps with 500 replications. Figure A3 in Web Appendix §D shows the bootstrapped distribution of these ATT estimates. The results in Table 8 suggest that consumers become more price-sensitive in the offline channel after they adopt online shopping, and this finding is consistent across these categories. The effect is stronger for the dog and cat hygiene categories, consistent with the elasticity visualization in Figure 2. To help with interpretation, we calculate the relative magnitude of the elasticity increase in percentage terms by dividing the ATT by the average pre-adoption price elasticity of adopters and show these percentage changes in the second row of Table 8. The most substantial effects are observed for dog hygiene and cat hygiene categories; offline price elasticity increases by 72.7% (i.e., -1.792 / -2.464) for dog hygiene and 34.7% (i.e., -0.738 / -2.125) for cat hygiene. The increase in offline price elasticity is lower for dry dog food, at 8.4% (i.e., -0.248 / -2.935). As discussed in §7.1, the pronounced increase in offline price elasticity for hygiene products (vs. dry food products) is likely driven by the lower switching costs associated with trying new brands of hygiene mats or other hygiene-related items. In sum, consistent with the descriptive evidence earlier, we find that consumers become more price-sensitive in their offline shopping trips when they start shopping online.
@@ 8.1 Generalizability
[…]
@@ 8.2 Robustness checks
We first examine alternative demand specifications in $\S$8.2.1 and then consider matching methods and alternative control groups in $\S$8.2.2 and 8.2.3, respectively. Finally, in $\S$8.2.4, we consider an alternative definition of the purchase occasions.
@@ 8.2.1 Alternative Demand Specifications
    Accounting for a potential change in state dependence or brand loyalty due to online adoption: In our main model, we include customer-brand fixed effects to capture time-invariant preferences and use $\mathcal{I}\{\textrm{BuyPrevOcca}_{ijt}\}$ to capture the state dependence. We now consider a specification with an additional control variable, $\mathcal{I}\{\textrm{BuyPrevOcca}_{ijt}\}\times\textrm{cumPercOnlineSpend}_{it}$, which captures the interaction between state dependence (e.g., repeated purchase of a brand) and the share of online spending. This can account for any effects that online adoption may have on brand loyalty.
[…]
@@ 8.2.2 Matching and IPW Approaches to Control for Selection on Observables
[…]
@@
We perform the propensity score estimation and matching for each adoption cohort separately. Specifically, we define the pre-adoption period separately for each cohort and measure pre-adoption variables (e.g., monthly average spend) using all purchase data up to the month of adoption (Gu and Kannan, 2021). For example, when identifying matched non-adopters for customers who adopt in July 2019, we measure the pre-adoption variables for both adopters and non-adopters using the data up to June 30, 2019, and estimate the propensity score model for the adoption cohort in July 2019. We repeat this process for each cohort and estimate twelve different propensity score models corresponding to each monthly adoption cohort. We compute propensity scores with a logistic regression model and the variables included are demographics (e.g., gender, age, and household income) and pre-adoption purchase variables (i.e. average monthly spend, number of unique orders, unique brands, and unique categories). With the estimated propensity scores, we use our CS estimator to estimate the ATT using both IPW and PSM methods. The propensity score model specification and detailed results from the IPW and PSM methods are shown in Web Appendix E.4. We find that our conclusions remain the same.